OpenStudy (anonymous):

If a 14-foot diameter sheave is used to drive a cable car, at what angular velocity must the sheave turn in order for the cable car to travel 11 miles per hour? Give your answer in revolution per minute. (1 mi = 5280ft)

OpenStudy (anonymous):

\[\frac{\text {11*}5280}{60 (\pi 14)}=\frac{22.0089 \text{ rev}}{\text{minute}}\]